Assuming an interest rate of 10 percent compounded semiannually, the future value of $500 in 16 years is . (Do not include the dollar sign ($). Round your answer to 2 decimal places. (e.g. 32.16)) References eBook& Resources
Answer #1

I= 10%÷2 = 5; principal = 500 ; n = 2× 16 =32

Future value = p * (1+I) ^n

= 500* (1+5%)^32

= 2382.47